# PersonalAI: A Systematic Comparison of Knowledge Graph Storage and Retrieval Approaches for Personalized LLM agents

### v2

@@ 8 Future work
Appendix A LLM prompts used to build a knowledge graph within the Memorize pipeline
@@
Tables 5 and 6 present the LLM prompts employed in the Memorize pipeline for extracting thesis-oriented and simple memory triples, respectively, from unstructured natural language text. These prompts facilitate the transformation of textual data into a structured knowledge graph representation.
@@
Table 5: LLM prompts for extracting thesis memories (in the form of triplets) from natural language text
@@
Table 6: LLM prompts for extracting simple memories (in the form of triplets) from natural language text
Appendix B LLM prompts used to find outdated information in the constructed knowledge graph
Tables 7 and 8 displays the LLM prompts employed in the Memorize pipeline for identifying stale thesis-related and simple memories, respectively, within the knowledge graph.
@@ Appendix B LLM prompts used to find outdated information in the constructed knowledge graph
Table 7: LLM prompts for identifying outdated thesis memories
@@
Table 8: LLM prompts for detecting obsolete simple memories
Appendix C LLM prompts used within proposed QA pipeline
Table 9 presents the LLM prompts employed in the QA pipeline at the second stage (the “Query LLM Parser” module) for extracting key entities from the original user question. Table 10 displays the LLM prompts used in the fourth stage (the “QA LLM Generator” module) to generate contextually appropriate responses to user queries.
@@ Appendix C LLM prompts used within proposed QA pipeline
Table 9: LLM prompts for extracting key entities from natural language text
@@
Table 10: LLM prompts for conditional generation of answer to user–question
Appendix D LLM–as–a–Judge instructions
To ensure the reproducibility of the obtained results, LLM inference was conducted using a deterministic generation strategy. The following hyperparameters were applied: num_predict – 2048, seed – 42, temperature – 0.0, and top_k – 1. The Qwen2.5 7B model, sourced from the Ollama repository, was prompted to evaluate whether the outputs of the proposed QA pipeline correctly answered the given questions. The specific LLM prompts used for this assessment are provided in Table 11.
Type Prompt
System [Uncaptioned image]
User [Uncaptioned image]
Assistant [Uncaptioned image]
Table 11: LLM prompts used in LLM–as–a–Judge framework
Appendix E Used preprocessing methods for evaluation datasets
For the original HotpotQA dataset, the distractor/validation subset was selected, comprising 7405 question-answer (QA) pairs and 13781 unique contexts. QA pairs were then filtered to exclude those with associated contexts falling outside a specified length range (measured in characters), retaining only contexts between 64 and 1024 characters in length. This filtering process resulted in 13291 remaining contexts. Finally, the first 2000 QA pairs and their corresponding contexts were extracted, yielding a final subset of 3933 unique contexts.
@@ Appendix E Used preprocessing methods for evaluation datasets
Thus, evaluation sets for quality assessment of the proposed/implemented Memorize– and QA–pipelines were obtained. The characteristics of the obtained subsets of HotpotQA–, TriviaQA– and DiaASQ–datasets can be found in Table 12.
Dataset QA–pairs Relevant contexts
@@
Table 12: Characteristics of the datasets used to evaluate the quality of the proposed Memorize– and QA–pipelines
Appendix F Retrieval hyperparameters
@@
Appendix G Characteristics of constructed knowledge graphs
@@
To evaluate the QA pipeline within the ablation experiments, we constructed 14 knowledge graphs based on the given dataset and LLM configurations. The structural characteristics of the generated graphs are detailed in Table 13.
@@
The results indicate that certain graphs experienced parsing errors during the extraction of LLM responses, resulting in incomplete storage of contextual information. Across the evaluated datasets, the average parsing error rates were as follows: DiaASQ (7.0%), HotpotQA (6.3%), and TriviaQA (7.3%). For the DiaASQ dataset, DeepSeekR1 7B produced the highest number of thesis and object memory nodes, as well as hyper-relations, while GPT4o-mini extracted the largest number of unique simple-relations. In contrast, for the HotpotQA and TriviaQA datasets, Qwen2.5 7B generated the most thesis and object memory nodes, whereas DeepSeekR1 7B again yielded the highest number of hyper-relations.
Based on the obtained results, we can model the structural characteristics of the knowledge graphs generated by each evaluated system: see Figure 14.
@@
The data reveal significant variation in LLM response parsing accuracy across models. Specifically, Qwen2.5 7B and Llama3.1 8B demonstrate the lowest error rates (0.02% each), while DeepSeek V3 exhibits the highest parsing error rate (31.21%). Intermediate performance is observed in DeepSeekR1 7B (0.29%) and GPT4o mini (9.87%). Regarding knowledge graph composition, DeepSeekR1 7B and Qwen2.5 7B yield the most comprehensive representations, generating the highest number of thesis/object memories and associated relations (both hyper-relational and simple). Further analysis of node creation efficiency per contextual unit shows Qwen2.5 7B achieves superior granularity, producing the largest number of thesis/object nodes while maintaining contextual coherence: see Table 15.
@@
LLM Number of contexts to store in graph Nubmer of nodes Number of relations Average number of matching vertices (by type) episodic thesis object episodic hyper simple object–neighbours (to episodic–nodes) thesis–neighbours (to episodic–nodes) object–neighbours (to thesis–nodes) object–neighbours (to object–nodes) DiaASQ DeepSeek R1 7B 3483 3477 34039 30097 138584 133301 34049 30.06 10.01 3.91 1.99 Qwen2.5 7B 3482 32512 28420 129974 111618 31290 27.98 9.48 3.43 2.04 Llama3.1 8B 3482 29655 20014 100045 77467 28063 20.21 8.52 2.61 2.25 GPT4o mini 3482 31361 28172 125403 96667 39477 27.00 9.00 3.08 1.89 DeepSeek V3 2270 21482 18416 80118 64349 28877 25.83 9,46 2.99 2.11 HotpotQA DeepSeek R1 7B 3933 3921 27137 55254 121359 111174 33714 24.00 7.34 4.09 1.55 Qwen2.5 7B 3933 31795 56078 119387 106037 38460 22.26 8.15 3.33 1.36 Llama3.1 8B 3933 26364 40601 89506 75332 29021 16.04 6.79 2.85 1.47 GPT4o mini 3933 30777 48771 105524 93791 42599 18.99 7.94 3.04 1.36 DeepSeek V3 2713 20164 35291 73242 63921 34621 19.55 7.53 3.17 1.41 TriviaQA DeepSeek R1 7B 4925 4905 48855 96132 213481 201861 52019 33.50 10.56 4.13 1.53 Qwen2.5 7B 4923 52835 109900 220991 188780 62043 34.15 10.87 3.57 1.27 Llama3.1 8B 4922 45241 72285 158202 127389 46757 22.83 9.35 2.81 1.54 DeepSeek V3 3506 37496 68602 143933 122480 61821 30.32 10.75 3.26 1.42
Table 13: Characteristics of the constructed knowledge graphs within conducted experiments
@@
LLM Number of contexts to store in graph Number of nodes Number or relations Average number of adjacent vertices (by type) episodic thesis object episodic hyper simple object–neighbours (to episodic–nodes) thesis–neighbours (to episodic–nodes) object–neighbours (to thesis–nodes) object–neighbours (to object–nodes) DeepSeek R1 7B 4113 4101 36677 60494 157808 148778 39927 29.18 9.30 4.04 1.69 Qwen2.5 7B 4112 39047 64799 156784 135478 43931 28.13 9.5 3.44 1.55 Llama3.1 8B 4112 33753 44300 115917 93396 34613 19.69 8.22 2.75 1.75 GPT4o mini 3707 31069 38471 115463 95229 41038 22.99 8.47 3.06 1.625 DeepSeek V3 2829 26380 40769 99097 83583 41773 25.23 9.24 3.14 1.64
@@
Table 14: Average/expected characteristics of knowledge graphs in case of using given LLM models for their construction
LLM Node Relation
thesis object hyper simple
@@
Table 15: The number of unique nodes/relations that are added to the knowledge graph when processing (using a given LLM model) and storing one episodic memory (episodic node)
@@ Appendix H Comparison with existing RAG and GraphRAG methods
To evaluate the performance of optimal configurations derived from our framework against existing Retrieval-Augmented Generation (RAG) and GraphRAG approaches, we conducted a systematic literature review. The search was performed across five academic search engines: SciSpace, Scite, PaperDigest, Consensus, and Elicit. Using each engine, we selected the first ten publications in the search results for the following queries: in the case of RAG methods – “Retrieval-augmented generation methods based on pretrained language models” and “RAG methods in NLP”; in the case of GraphRAG methods – “RAG on Knowledge Graphs”, “Enhancing RAG–approach with Knowledge Graphs”, “Graph RAG” and “RAG with integration of Large Language Models (LLMs) and Knowledge Graphs (KGs)”. The search was restricted to publications up to 2018. Subsequently, we applied a three-stage filtering process: (1) duplicate entries were removed, and papers introducing novel evaluation datasets were excluded; (2) works focusing on domain-specific applications without broader methodological contributions were discarded; (3) only studies providing comprehensive methodological descriptions and employing standardized benchmarks were retained. This process yielded two final sets of nine articles each, covering RAG and GraphRAG techniques, respectively. A comparative analysis of these methods against our framework’s optimal graph construction and retrieval configuration is presented in Table 16, with performance measured using ExactMatch metric.
Dataset RAG–method Our method
REALM DPR RAG ColBERT–QA FiD EMDR2 RETRO Atlas RePLUG
@@
GraphRAG–method
ToG RoG PMKGE GRAG GNN–RAG ToG2.0 DoG GCR PDA
@@
Table 16: Comparison of existing RAG– and GraphRAG–methods with our proposed method on ExactMatch–metric. On TriviaQA dataset QA–configuration with DeepSeek V3 and combination of BeamSearch and WaterCircles algorithms (without restrictions on graph traversal) were used. On HotpotQA dataset QA–configuration with GPT4o–mini and the same algorithm and graph restrictions were used.
As shown in Table 16, our proposed method achieves a 14.1% improvement over existing GraphRAG approaches in a specific configuration. However, it underperforms compared to standard RAG methods by 17.8%. This discrepancy can be attributed to the fact that the RAG baselines evaluated in this study employed Reader and Retriever models that were specifically fine-tuned on the same dataset used for evaluation. As established in prior works, such in-domain fine-tuning typically yields optimal performance, and evaluation on out-of-domain datasets would be expected to result in significant degradation of measured metrics.

### v3

@@ VIII Future work
Appendix A LLM prompts used to build memory graph by Memorize pipeline
@@
Tables VII and VIII presents LLM prompts employed in Memorize pipeline for extracting thesis and object memory triples, respectively, from unstructured natural language text. These prompts facilitate the transformation of textual data into a structured knowledge graph representation.
@@
TABLE VII: LLM prompts for extracting thesis memories (in the form of triples) from natural language text
@@
TABLE VIII: LLM prompts for extracting object memories (in the form of triples) from natural language text
Appendix B LLM prompts used to find outdated information in constructed memory graph
Tables IX and X displays LLM prompts employed in Memorize pipeline for identifying stale thesis-related and object memories, respectively, in memory graph.
@@ Appendix B LLM prompts used to find outdated information in constructed memory graph
TABLE IX: LLM prompts for identifying outdated thesis memories
@@
TABLE X: LLM prompts for detecting obsolete object memories
Appendix C LLM prompts used in proposed QA pipeline
Table XI presents LLM prompts employed in QA pipeline at the second stage for extracting key entities from the original user question. Table XII displays LLM prompts used in the fourth stage to generate contextually appropriate responses to user question.
@@ Appendix C LLM prompts used in proposed QA pipeline
TABLE XI: LLM prompts for extracting key entities from natural language text
@@
TABLE XII: LLM prompts for conditional generation of answer to user question
Appendix D Preprocessing operations for evaluation datasets
For the original HotpotQA dataset, the distractor/validation subset was selected, comprising 7405 question-answer (QA) pairs and 13781 unique contexts. QA pairs were then filtered to exclude those with associated contexts falling outside a specified length range (measured in characters), retaining only contexts between 64 and 1024 characters in length. This filtering process resulted in 13291 remaining contexts. Finally, first 2000 QA pairs and their corresponding contexts were extracted, yielding a final subset of 3933 unique contexts.
@@ Appendix D Preprocessing operations for evaluation datasets
Thus, evaluation sets for evaluation of proposed/implemented Memorize and QA pipelines were obtained. The characteristics of the obtained subsets of HotpotQA, TriviaQA and DiaASQ datasets can be found in Table XIII.
Dataset QA-pairs Relevant contexts
@@
TABLE XIII: Extended characteristics of datasets, used to evaluate proposed Memorize and QA pipelines
Appendix E Retrieval hyperparameters
@@
Appendix F LLM–as–a–Judge instructions
To ensure the reproducibility of the obtained results, LLM inference was conducted using a deterministic generation strategy. The following hyperparameters were applied: num_predict – 2048, seed – 42, temperature – 0.0, and top_k – 1. The Qwen2.5 7B model, sourced from the Ollama repository, was prompted to evaluate whether the outputs of the proposed QA pipeline correctly answered the given questions. The specific LLM prompts used for this assessment are provided in Table XIV.
Type Prompt
System [Uncaptioned image]
User [Uncaptioned image]
Assistant [Uncaptioned image]
TABLE XIV: LLM prompts used in LLM–as–a–Judge framework
Appendix G Characteristics of constructed memory graphs
@@
To evaluate the QA pipeline we constructed 14 memory graphs based on the given dataset and LLM configurations. The structural characteristics of the generated graphs are detailed in Table XV.
@@
Results indicate that certain graphs experienced parsing errors during the extraction of LLM responses, resulting in incomplete storage of contextual information. Across the evaluated datasets, the average parsing error rates were as follows: DiaASQ (7.0%), HotpotQA (6.3%), and TriviaQA (7.3%). For the DiaASQ dataset, DeepSeekR1 7B produced highest number of thesis and object vertices, as well as hyper-edges, while GPT4o-mini extracted largest number of unique object-typed edges. In contrast, for the HotpotQA and TriviaQA datasets, Qwen2.5 7B generated most thesis and object vertices, whereas DeepSeekR1 7B again yielded the highest number of hyper-edges. Based on the obtained results, we can model the structural characteristics of memory graphs generated by each evaluated system: see Figure XVI.
@@
The data reveal significant variation in LLM response parsing accuracy across models. Specifically, Qwen2.5 7B and Llama3.1 8B demonstrate lowest error rates (0.02% each), while DeepSeek V3 exhibits the highest parsing error rate (31.21%). Intermediate performance is observed in DeepSeekR1 7B (0.29%) and GPT4o mini (9.87%). Regarding memory graph composition, DeepSeekR1 7B and Qwen2.5 7B yield the most comprehensive representations, generating the highest number of thesis/object memories and associated edges. Further analysis of vertices creation efficiency per contextual unit shows Qwen2.5 7B achieves superior granularity, producing the largest number of thesis/object memories while maintaining contextual coherence: see Table XVII.
@@
LLM Number of contexts to store in graph Nubmer of vertices Number of edges Average number of matching vertices (by type) episodic thesis object episodic thesis object object neighbours (to episodic vertices) thesis neighbours (to episodic vertices) object neighbours (to thesis vertices) object neighbours (to object vertices) DiaASQ DeepSeek R1 7B 3483 3477 34039 30097 138584 133301 34049 30.06 10.01 3.91 1.99 Qwen2.5 7B 3482 32512 28420 129974 111618 31290 27.98 9.48 3.43 2.04 Llama3.1 8B 3482 29655 20014 100045 77467 28063 20.21 8.52 2.61 2.25 GPT4o mini 3482 31361 28172 125403 96667 39477 27.00 9.00 3.08 1.89 DeepSeek V3 2270 21482 18416 80118 64349 28877 25.83 9,46 2.99 2.11 HotpotQA DeepSeek R1 7B 3933 3921 27137 55254 121359 111174 33714 24.00 7.34 4.09 1.55 Qwen2.5 7B 3933 31795 56078 119387 106037 38460 22.26 8.15 3.33 1.36 Llama3.1 8B 3933 26364 40601 89506 75332 29021 16.04 6.79 2.85 1.47 GPT4o mini 3933 30777 48771 105524 93791 42599 18.99 7.94 3.04 1.36 DeepSeek V3 2713 20164 35291 73242 63921 34621 19.55 7.53 3.17 1.41 TriviaQA DeepSeek R1 7B 4925 4905 48855 96132 213481 201861 52019 33.50 10.56 4.13 1.53 Qwen2.5 7B 4923 52835 109900 220991 188780 62043 34.15 10.87 3.57 1.27 Llama3.1 8B 4922 45241 72285 158202 127389 46757 22.83 9.35 2.81 1.54 DeepSeek V3 3506 37496 68602 143933 122480 61821 30.32 10.75 3.26 1.42
TABLE XV: Characteristics of constructed memory graphs for QA experiments
@@
LLM Number of contexts to store in graph Number of vertices Number of edges Average number of adjacent vertices (by type) episodic thesis object episodic thesis object object neighbours (to episodic vertices) thesis neighbours (to episodic vertices) object neighbours (to thesis vertices) object neighbours (to object vertices) DeepSeek R1 7B 4113 4101 36677 60494 157808 148778 39927 29.18 9.30 4.04 1.69 Qwen2.5 7B 4112 39047 64799 156784 135478 43931 28.13 9.5 3.44 1.55 Llama3.1 8B 4112 33753 44300 115917 93396 34613 19.69 8.22 2.75 1.75 GPT4o mini 3707 31069 38471 115463 95229 41038 22.99 8.47 3.06 1.625 DeepSeek V3 2829 26380 40769 99097 83583 41773 25.23 9.24 3.14 1.64
@@
TABLE XVI: Average/expected characteristics of knowledge graphs in case of using given LLM models for their construction
LLM Vertex Edge
thesis object thesis object
@@
TABLE XVII: Number of unique vertices/edges that were added to the graph when processing (using a given LLM model) and storing one episodic memory (episodic verrtex)
In addition to characteristics of constructed graphs, we collected information about time and speed of the Memorize pipeline, which responsible for parsing incoming unstructured texts in natural language and storing them in memory model: see Table XVIII.
LLM / Dataset Memory construction time (hours) and speed (contexts per minute) Mean:
Qwen2.5 7B DeepSeek R1 7B Llama3.1 8B GPT4o mini DeepSeek V3
DiaASQ 23.5 / 2.47 57.5 / 1.00 32 / 1.81 32 / 1.81 68 / 0.85 42.6 / 1.58
HotpotQA 61.5 / 1.06 47.5 / 1.38 27 / 2.42 34 / 1.92 72 / 0.91 48.4 / 1.53
TriviaQA 90 / 0.91 - 90 / 0.91 - 80 / 1.02 86.6 / 0.94
Mean: 58.3 / 1.48 52.5 / 1.19 49.7 / 1.71 33 / 1.86 73.3 / 0.92
TABLE XVIII: Time and speed of memory graphs construction based on a given dataset and LLM, which were used in proposed QA pipeline evaluation
Table XVIII shows that highest speed of parsing and saving information into memory is observed in configurations of Memorize pipeline there GPT4o mini is set as LLM (1.86); in second place is LLama3.1 8B, that was deployed in Ollama Docker container on local machine.
It is also important to note the amount of disk space that constructed memory is required. When using Milvus as a database for storing vector representations of text, on selected datasets, one memory graph in our implementation occupies, approximately, 80-90 GB. If Qdrant is used as a vector storage, the same graphs will occupy, approximately, 4-6 GB each.
@@ Appendix H Comparison with existing RAG and GraphRAG methods
Based on experimental results, a comparative table, summarizing best-performing QA configurations by Exact Match metric, was compiled (see Table XIX).
LLM /
Dataset
Qwen2.5 7B DeepSeek R1 7B Llama3.1 8B GPT4o mini DeepSeek V3
DiaASQ 0.22 / BS + WC / T 0.11 / AS / E 0.18 / AS / E 0.47 / BS + WC / E 0.46 / BS + WC / O
HotpotQA 0.18 / AS / all 0.14 / AS / E 0.37 / BS / O 0.59 / BS + WC / all 0.6 / BS + WC / T
TriviaQA 0.2 / BS / E 0.18 / AS / E 0.47 / BS / E - 0.62 / BS + WC / all
Mean: 0.2 0.14 0.34 0.53 0.56
TABLE XIX: Best QA configurations ranked by Exact Match metric across all experiments. The corresponding cells contains Exact Match score, used retrieval algorithm and the type of restriction applied to the memory graph during traversal. Shortcuts for retrieval algorithms: BA – BeamSearch; AS – A*; BS + AS – hybrid of BeamSearch and A*; BS + WC – hybrid of BeamSearch and WaterCircles. Shortcuts for graph restrictions: all – no restrictions applied; E – episodic vertices excluded from traversal; T – thesis vertices excluded; O – object vertices excluded.
As shown in Table XIX, Qwen2.5 achieve best performance (0.2) among the 7B models. Among all evaluated configurations, highest overall effectiveness (0.56) was reached by setups incorporating DeepSeek V3. Notably, top-performing 7B configurations predominantly relied on A*, especially under constraint that traversal restricted through episodic vertices. In contrast, best DeepSeek V3 configurations frequently adopted a hybrid strategy combining BeamSearch and WaterCircles. Across high-performing configurations more broadly, BeamSearch consistently appeared as a key component within the retrieval pipeline.
To evaluate the performance of optimal configurations derived from our framework against existing Retrieval-Augmented Generation (RAG) and GraphRAG approaches, we conducted a systematic literature review. The search was performed across five academic search engines: SciSpace, Scite, PaperDigest, Consensus, and Elicit. Using each engine, we selected the first ten publications in the search results for the following queries: in case of RAG methods – ”Retrieval-augmented generation methods based on pretrained language models” and ”RAG methods in NLP”; in case of GraphRAG methods – ”RAG on Knowledge Graphs”, ”Enhancing RAG–approach with Knowledge Graphs”, ”Graph RAG” and ”RAG with integration of Large Language Models (LLMs) and Knowledge Graphs (KGs)”. The search was restricted to publications up to 2018. Subsequently, we applied a three-stage filtering process: (1) duplicate entries were removed, and papers introducing novel evaluation datasets were excluded; (2) works focusing on domain-specific applications without broader methodological contributions were discarded; (3) only studies providing comprehensive methodological descriptions and employing standardized benchmarks were retained. This process yielded two final sets of nine articles each, covering RAG and GraphRAG techniques, respectively. A comparative analysis of these methods against our framework’s optimal graph construction and retrieval configuration is presented in Table XX, with performance measured using ExactMatch metric.
Dataset RAG method Our method
REALM [10.5555/3524938.3525306] DPR [karpukhin-etal-2020-dense] RAG [10.5555/3495724.3496517] ColBERT–QA [khattab-etal-2021-relevance] FiD [izacard-grave-2021-leveraging] EMDR2 [sachan2021endtoend] RETRO [Borgeaud2021ImprovingLM] Atlas [10.5555/3648699.3648950] RePLUG [shi-etal-2024-replug]
@@
GraphRAG method
ToG [sun2023thinkongraph] RoG [luo2024rog] PMKGE [liu2025enhancinglargelanguagemodels] GRAG [hu2024grag] GNN–RAG [mavromatis-karypis-2025-gnn] ToG2.0 [Ma2024ThinkonGraph2D] DoG [ma2025debate] GCR [luo2024graph] PDA [Sun2024PyramidDrivenAP]
@@
TABLE $X$: Comparison of existing RAG and GraphRAG methods with our proposed method on ExactMatch metric. On TriviaQA dataset QA configuration with DeepSeek V3 and combination of BeamSearch and WaterCircles algorithms (without restrictions on graph traversal) were used. On HotpotQA dataset QA configuration with GPT4o–mini and the same algorithm and graph restrictions were used.
As shown in Table XX, our proposed method achieves a 14.1% improvement over existing GraphRAG approaches in a specific configuration. However, it underperforms compared to standard RAG methods by 17.8%. This discrepancy can be attributed to the fact that the RAG baselines evaluated in this study employed Reader and Retriever models that were specifically fine-tuned on the same dataset used for evaluation. As established in prior works, such in-domain fine-tuning typically yields optimal performance, and evaluation on out-of-domain datasets would be expected to result in significant degradation of measured metrics.